# Functional Matching of Logic Subgraphs: Beyond Structural Isomorphism

### v1

@@ Abstract
Subgraph matching in logic circuits is foundational for numerous Electronic Design Automation (EDA) applications, including datapath optimization, arithmetic verification, and hardware trojan detection. However, existing techniques rely primarily on structural graph isomorphism and thus fail to identify function-related subgraphs when synthesis transformations substantially alter circuit topology. To overcome this critical limitation, we introduce the concept of functional subgraph matching, a novel approach that identifies whether a given logic function is implicitly present within a larger circuit, irrespective of structural variations induced by synthesis or technology mapping. Specifically, we propose a two-stage multi-modal framework: (1) learning robust functional embeddings across AIG and post-mapping netlists for functional subgraph detection, and (2) identifying fuzzy boundaries using a graph segmentation approach. Evaluations on standard benchmarks (ITC99, OpenABCD, ForgeEDA) demonstrate significant performance improvements over existing structural methods, with average $93.8\%$ accuracy in functional subgraph detection and a dice score of $91.3\%$ in fuzzy boundary identification.
@@ 2.3 Functional Subgraph
  • Reflexivity: For any graph $\mathcal{G}$, $\mathcal{G}$ is the functional subgraph of $\mathcal{G}$, i.e. $\forall\mathcal{G},\mathcal{G}\preccurlyeq\mathcal{G}$
  • Functional Equivalence Preservation: If $\mathcal{G}_{1}$ is a functional subgraph of $\mathcal{G}_{2}$, and $\mathcal{G}^{\prime}_{1}$ is functionally equivalent to $\mathcal{G}_{1}$, then $\mathcal{G}^{\prime}_{1}$ is a functional subgraph of $\mathcal{G}_{2}$, i.e. if $\mathcal{G}_{1}\preccurlyeq\mathcal{G}_{2}$ and $\mathcal{G}^{\prime}_{1}\equiv_{func}\mathcal{G}_{1}$, then $\mathcal{G}^{\prime}_{1}\preccurlyeq\mathcal{G}_{2}$.
@@ 3.1 Stage #1: Functional Subgraph Detection
EDA flows such as synthesis and mapping modify the circuit structure while preserving functional equivalence. As defined in Definition 3, the functional subgraph relation focuses on the functionality of the candidate circuits rather than structure, as they can be transformed into an equivalent circuit with any structure. Therefore, learning function-invariant embeddings for equivalent circuits across different stages is crucial for functional subgraph detection. While $\mathcal{G}_{aig}$ and $\mathcal{G}_{syn}$ share the same gate types, $\mathcal{G}_{aig}$ and $\mathcal{G}_{pm}$ differ significantly in modality, i.e., the gate types in $\mathcal{G}_{pm}$ are substantially dissimilar to those in $\mathcal{G}_{aig}$. Therefore, we employ both intra-modal and inter-modal alignment techniques to acquire function-invariant and structure-agnostic embeddings with the InfoNCE loss [33]. We select $\mathcal{G}_{aig}$ as the anchor and compute the intra-modal and inter-modal losses as follows:
@@ 3.2 Stage #2: Fuzzy Boundary Identification
Given a post-mapping netlist $\mathcal{G}_{pm}$, we replace the cells in $\mathcal{G}_{pm}$ with the corresponding AIGs from the library to acquire the netlist $\mathcal{G}_{aig}$. This process yields a mapping function $\phi$ that associates the node indices of $\mathcal{G}_{aig}$ with those of $\mathcal{G}_{pm}$. Next, we randomly drop nodes to obtain $\mathcal{G}_{sub}$, which serves as the functional subgraph of both $\mathcal{G}_{pm}$ and $\mathcal{G}_{aig}$. Using the subgraph $\mathcal{G}_{sub}$, we annotate the nodes in $\mathcal{G}_{pm}$ by mapping the node indices of $\mathcal{G}_{sub}$ to those of $\mathcal{G}_{pm}$ through the function $\phi$. Specifically, for each node in $\mathcal{G}_{sub}$, if it maps to a node $i$ in $\mathcal{G}_{pm}$, we annotate node $i$ as 1; otherwise, we annotate it as 0. This annotation process strictly follows the fuzzy boundary definition in Definition 4.
@@ 4.3 Stage #2: Fuzzy Boundary Identification
[…]
@@ Appendix A Proofs of the Proposed Properties
If $\mathcal{G}_{1}\preccurlyeq\mathcal{G}_{2}$ and $\mathcal{G}^{\prime}_{1}\equiv_{func}\mathcal{G}_{1}$, then $\mathcal{G}^{\prime}_{1}\preccurlyeq\mathcal{G}_{2}$.
@@
Proposition 6 (Anti-symmetry).
@@ B.2 Fuzzy Boundary Identification
We futher evalute these methods on fuzzy boundary identification. The results are shown in Table 5.
@@ Appendix C Datasets and Implementation Details
Dataset statistics and splits are shown in Table  6. For dataset split, we first split the training circuits and test circuits in the source dataset, then we cut subgraph for the training circuit and test circuits to generate our small circuit dataset. For ITC99 and OpenABCD, the split follow the previous work [24]. For ForgeEDA, we randomly select $10\%$ circuits in the dataset as test circuits. For small circuit, we apply Algorithm 1 to randomly sample subgraph.
Source Dataset Split #Pair $\mathcal{G}_{sub}$ $\mathcal{G}_{aig}$ $\mathcal{G}_{syn}$ $\mathcal{G}_{pm}$
#Node Depth #Node Depth #Node Depth #Node Depth
Environment
Evaluation Metrics
[…]

### v2

@@ Abstract
Subgraph matching in logic circuits is foundational for numerous Electronic Design Automation (EDA) applications, including datapath optimization, arithmetic verification, and hardware trojan detection. However, existing techniques rely primarily on structural graph isomorphism and thus fail to identify function-related subgraphs when synthesis transformations substantially alter circuit topology. To overcome this critical limitation, we introduce the concept of functional subgraph matching, a novel approach that identifies whether a given logic function is implicitly present within a larger circuit, irrespective of structural variations induced by synthesis or technology mapping. Specifically, we propose a two-stage multi-modal framework: (1) learning robust functional embeddings across AIG and post-mapping netlists for functional subgraph detection, and (2) identifying fuzzy boundaries using a graph segmentation approach. Evaluations on standard benchmarks (ITC99, OpenABCD, ForgeEDA) demonstrate significant performance improvements over existing structural methods, with average $93.8\%$ accuracy in functional subgraph detection and a dice score of $91.3\%$ in fuzzy boundary identification. The source code and implementation details can be found at our repository.
@@ 2.3 Functional Subgraph
  • Reflexivity: For any graph $\mathcal{G}$, $\mathcal{G}$ is the functional subgraph of $\mathcal{G}$, i.e. $\forall\mathcal{G},\mathcal{G}\preccurlyeq\mathcal{G}$.
  • Functional Equivalence Preservation: If $\mathcal{G}_{1}$ is a functional subgraph of $\mathcal{G}_{2}$, then:
    • (Left-hand Side) if $\mathcal{G}^{\prime}_{1}$ is functionally equivalent to $\mathcal{G}_{1}$, then $\mathcal{G}^{\prime}_{1}$ is a functional subgraph of $\mathcal{G}_{2}$, i.e. if $\mathcal{G}_{1}\preccurlyeq\mathcal{G}_{2}$ and $\mathcal{G}^{\prime}_{1}\equiv_{func}\mathcal{G}_{1}$, then $\mathcal{G}^{\prime}_{1}\preccurlyeq\mathcal{G}_{2}$.
    • (Right-hand Side) if $\mathcal{G}^{\prime}_{2}$ is functionally equivalent to $\mathcal{G}_{2}$, then $\mathcal{G}_{1}$ is a functional subgraph of $\mathcal{G}^{\prime}_{2}$, i.e. if $\mathcal{G}_{1}\preccurlyeq\mathcal{G}_{2}$ and $\mathcal{G}^{\prime}_{2}\equiv_{func}\mathcal{G}_{2}$, then $\mathcal{G}_{1}\preccurlyeq\mathcal{G}^{\prime}_{2}$.
@@ 3.1 Stage #1: Functional Subgraph Detection
EDA flows such as synthesis and mapping modify the circuit structure while preserving functional equivalence. As defined in Definition 3, the functional subgraph relation focuses on the functionality of the candidate circuits rather than structure, as they can be transformed into an equivalent circuit with any structure. Furthermore, the Functional Equivalence Preservation property in Section 2.3 imply that, if the subgraph relation $Q\preccurlyeq G$ hold, then if we replace $Q$ or $G$ with another functional equivalent graph, the subgraph relation continues to hold. This invariance is the key insight for our alignment: the embeddings of functionally equivalent graphs should be aligned, regardless of their structural variations.
Therefore, learning function-invariant embeddings for equivalent circuits across different stages is crucial for functional subgraph detection. While $\mathcal{G}_{aig}$ and $\mathcal{G}_{syn}$ share the same gate types, $\mathcal{G}_{aig}$ and $\mathcal{G}_{pm}$ differ significantly in modality, i.e., the gate types in $\mathcal{G}_{pm}$ are substantially dissimilar to those in $\mathcal{G}_{aig}$. Therefore, we employ both intra-modal and inter-modal alignment techniques to acquire function-invariant and structure-agnostic embeddings with the InfoNCE loss [33]. We select $\mathcal{G}_{aig}$ as the anchor and compute the intra-modal and inter-modal losses as follows:
@@ 3.2 Stage #2: Fuzzy Boundary Identification
As illustrated in Figure 3, given a post-mapping netlist $\mathcal{G}_{pm}$, we replace the cells in $\mathcal{G}_{pm}$ with the corresponding AIGs from the library to acquire the netlist $\mathcal{G}_{aig}$. This process yields a mapping function $\phi$ that associates the node indices of $\mathcal{G}_{aig}$ with those of $\mathcal{G}_{pm}$. Next, we randomly drop nodes to obtain $\mathcal{G}_{sub}$, which serves as the functional subgraph of both $\mathcal{G}_{pm}$ and $\mathcal{G}_{aig}$. Using the subgraph $\mathcal{G}_{sub}$, we annotate the nodes in $\mathcal{G}_{pm}$ by mapping the node indices of $\mathcal{G}_{sub}$ to those of $\mathcal{G}_{pm}$ through the function $\phi$. Specifically, for each node in $\mathcal{G}_{sub}$, if it maps to a node $i$ in $\mathcal{G}_{pm}$, we annotate node $i$ as 1; otherwise, we annotate it as 0. This annotation process strictly follows the fuzzy boundary definition in Definition 4.
@@ 4.3 Stage #2: Fuzzy Boundary Identification
4.4 Ablation Study
Table 3: Ablation Study on ITC99 Dataset($\%$).
@@ 6 Conclusion
Acknowledgment
This work was supported in part by the Hong Kong Research Grants Council (RGC) under Grant No. 14212422, 14202824, and C6003-24Y.
@@ Appendix A Proofs of the Proposed Properties
If $\mathcal{G}_{1}$ is a functional subgraph of $\mathcal{G}_{2}$, then:
  • (Left-hand Side) if $\mathcal{G}_{1}\preccurlyeq\mathcal{G}_{2}$ and $\mathcal{G}^{\prime}_{1}\equiv_{func}\mathcal{G}_{1}$, then $\mathcal{G}^{\prime}_{1}\preccurlyeq\mathcal{G}_{2}$.
  • (Right-hand Side) if $\mathcal{G}_{1}\preccurlyeq\mathcal{G}_{2}$ and $\mathcal{G}^{\prime}_{2}\equiv_{func}\mathcal{G}_{2}$, then $\mathcal{G}_{1}\preccurlyeq\mathcal{G}^{\prime}_{2}$.
@@
(Right-hand Side) According to the Definition 3, if $\mathcal{G}_{1}$ is a functional subgraph of $\mathcal{G}_{2}$, then there exist $\mathcal{G}^{*}_{2}$ that $\mathcal{G}^{*}_{2}\equiv_{\mathrm{func}}\mathcal{G}_{2}$ and $\mathcal{G}_{1}$ is an isomorphic subgraph of $\mathcal{G}^{*}_{2}$. Since $\mathcal{G}^{\prime}_{2}\equiv_{\mathrm{func}}\mathcal{G}_{2}$ and $\mathcal{G}^{*}_{2}\equiv_{\mathrm{func}}\mathcal{G}_{2}$, then $\mathcal{G}^{*}_{2}$. Since $\mathcal{G}^{\prime}_{2}\equiv_{\mathrm{func}}\mathcal{G}^{\prime}_{2}$. By the Definition 3, $\mathcal{G}_{1}$ is a functional subgraph of $\mathcal{G}^{\prime}_{2}$.
(Left-hand Side) By Definition 3, there exists a graph $\mathcal{G}_{2}^{\prime}\equiv_{\mathrm{func}}\mathcal{G}_{2}$, such that
@@
Proposition 6 (Anti‐symmetry).
@@ B.2 Fuzzy Boundary Identification
We further evaluate these methods on fuzzy boundary identification. The results are shown in Table 5.
@@
B.3 Scalability on Medium-Sized Circuits
we collect a medium-sized graph dataset from ForgeEDA [17], containing circuits with graph sizes ranging from 100 to 10000 nodes. The statistical information of the medium-sized dataset is shown in Table 6.
Table 6: Statistics of the medium-sized dataset.
Graph Type Nodes Edges Depth
$G_{sub}$ $192.1\pm 320.87$ $207.16\pm 354.01$ $27.91\pm 30.08$
$G_{syn}$ $1396.99\pm 1764.63$ $1958.84\pm 2519.23$ $66.48\pm 99.7$
$G_{pm}$ $679.93\pm 851.2$ $1352.96\pm 1703.16$ $21.83\pm 32.6$
We perform functional subgraph detection on this dataset, and the results are shown in Table 7 and  8. Since ABGNN encounters out-of-memory error when encoding graphs with deep logic levels, we do not report its results on this new dataset. While our method still demonstrates state-of-the-art performance, it shows a significant performance drop (from an F1-score of 95.3% to 81.2%, as shown in Table 1). This result highlights the challenge of scaling to larger circuits. We hope future work will explore and address this challenge.
Table 7: Functional Subgraph Detection on $G_{sub}\to G_{syn}$.
Method Accuracy Precision Recall F1-score
NeuroMatch $51.2_{\pm 3.3}$ $34.6_{\pm 24.5}$ $66.7_{\pm 47.1}$ $45.5_{\pm 32.2}$
HGCN $50.0_{\pm 0.0}$ $16.7_{\pm 23.6}$ $33.3_{\pm 47.1}$ $22.2_{\pm 31.4}$
Gamora $50.0_{\pm 0.0}$ $40.0_{\pm 14.1}$ $66.7_{\pm 47.1}$ $44.6_{\pm 31.3}$
Ours $\mathbf{81.5_{\pm 0.6}}$ $\mathbf{82.7_{\pm 1.1}}$ $\mathbf{79.8_{\pm 1.7}}$ $\mathbf{81.2_{\pm 0.8}}$
Table 8: Functional Subgraph Detection on $G_{sub}\to G_{pm}$.
Method Accuracy Precision Recall F1-score
NeuroMatch $51.0_{\pm 1.2}$ $33.9_{\pm 24.0}$ $66.6_{\pm 47.1}$ $44.9_{\pm 31.8}$
HGCN $50.0_{\pm 0.0}$ $16.7_{\pm 23.6}$ $33.3_{\pm 47.1}$ $22.2_{\pm 31.4}$
Gamora $50.0_{\pm 0.0}$ $33.3_{\pm 23.6}$ $66.7_{\pm 47.1}$ $44.4_{\pm 31.4}$
Ours $\mathbf{78.9_{\pm 1.0}}$ $\mathbf{80.6_{\pm 1.6}}$ $\mathbf{76.3_{\pm 2.7}}$ $\mathbf{78.3_{\pm 1.3}}$
B.4 Comparison with VF3
we evaluate the state-of-the-art subgraph isomorphic heuristic algorithm VF3 [34], which consistently achieves 100% precision and 100% recall on standard subgraph isomorphism tasks. Due to time constraints, we sampled circuits with fewer than 50 nodes from the ForgeEDA dataset and applied the VF3 algorithm. The results are shown in Table 9. According to our Definition 3 of Functional Subgraph, if $Q$ is an isomorphic subgraph of $G$, then $Q$ is always a functional subgraph of $G$. This is demonstrated by the 100% precision achieved by VF3. However, due to the function-preserving transformation, the explicit structure of $Q$ often disappears, leading to extremely low recall(0.32%) for VF3. These results highlight the importance of task definition.
Table 9: Comparison of subgraph isomorphism methods on different tasks.
$G_{sub}\to G_{syn}$ $G_{sub}\to G_{pm}$
Method Runtime (s) Precision Recall F1-score Precision Recall F1-score
VF3 480.8 100.0 0.32 0.65
Ours 8.0 88.5 91.9 90.2 86.4 94.5 90.3
@@ Appendix C Datasets and Implementation Details
Dataset statistics and splits are shown in Table  10. For dataset split, we first split the training circuits and test circuits in the source dataset, then we cut subgraph for the training circuit and test circuits to generate our small circuit dataset. For ITC99 and OpenABCD, the split follow the previous work [24]. For ForgeEDA, we randomly select $10\%$ circuits in the dataset as test circuits. For small circuit, we apply Algorithm 1 to randomly sample subgraph.